OpenStudy (anonymous):

For positive integers x, let ♮x be defined as the remainder when x is divided by 3. Which of the following is equal to ♮11? a. ♮16-♮5 b. ♮15-♮4 c. ♮14-♮3 d. ♮13-♮2 e. ♮12-♮1

OpenStudy (anonymous):

what kind of symbol is that?

OpenStudy (anonymous):

if you divide 11 by 3 the remainder is 2

OpenStudy (anonymous):

its a natural note symbol. the SAT question just had it ;_;

OpenStudy (anonymous):

so you have a bunch of work to do to see which of the above is equal 2

OpenStudy (anonymous):

for example \[ ♮16-♮5\] is the same as \(1-2=-1\)

OpenStudy (anonymous):

wait, why does ♮16 = 1

OpenStudy (anonymous):

which is in fact equal to 2 mod 3, but i am not exactly sure if that is what you want

OpenStudy (anonymous):

\(♮16 = 1\) because if you divide 16 by 3, the remainder is 1

OpenStudy (anonymous):

i.e. 3 goes in to 16 five times, with a remainder of 1

OpenStudy (anonymous):

so ♮x isn't the remainder?

OpenStudy (anonymous):

it is the remainder, yes

OpenStudy (anonymous):

that is why \( ♮16=1\)

OpenStudy (anonymous):

oh, okay. i think i get it xD

OpenStudy (anonymous):

the next one is \[ ♮15-♮4=0-1=-1\]

OpenStudy (anonymous):

ohhh.

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

\[ ♮14-♮3=2-0=2\] that is probably the one you want

OpenStudy (anonymous):

thx again. :]

OpenStudy (anonymous):

yw